# Preserving Expert-Level Privacy in Offline Reinforcement Learning

### v1

@@ 3.3 Problem Setup
Let $\Pi$ denote a set of $m$ private policies, corresponding to $m$ experts, $\Pi=\{\pi_{1},\dots,\pi_{m}\}$. $D^{\Pi}=\{D^{\pi_{1}},\dots,D^{\pi_{m}}\}$ is the aggregated non-private dataset generated by the interaction of the $m$ private policies with the environment. Each $D^{\pi_{i}}$ is a set of $N$ trajectories $\tau=(s_{1},a_{2},s_{2},a_{2},\dots,s_{L},a_{L},s_{L+1})$ of length $L$ obtained by logging the interactions of policy $\pi_{i}$ with the MDP $\mathcal{M}=\langle S,A,r,P,\rho_{0},\gamma\rangle$, where $s_{1}\sim\rho_{0}$, $a_{h}\sim\pi_{i}(\cdot|s_{h})$ and $s_{h+1}\sim P(\cdot|s_{h},a_{h})$. For notational simplicity we omit rewards throughout the paper as they do not play a role in our privatization approach and we can just assume they can be computed as needed for a given state-action pair.
We denote by $|\tau|$ the length of the trajectory as the number of full state-action pairs it consists of. Further, we use $\tau_{h:h^{\prime}}=(s_{h},a_{h},\dots s_{h},a_{h},s_{h^{\prime}+1})$ to denote the sub-trajectory of $\tau$ starting from the $h$-th timestep until the $h^{\prime}$-th, including the trailing next state $s_{h^{\prime}+1}$.
@@
The goal is to learn a policy using an offline RL algorithm with expert-level DP guarantees on $D^{\Pi}$, thus ensuring that the policy learnt in this manner is not much different from a policy learnt using $D^{\Pi^{\prime}}$. We make the following assumptions.
Assumption 1.
@@
This is a fairly mild assumption satisfied in many natural settings. The assumption arises in the first-phase of identifying stable trajectory prefixes for our algorithm, and extending this to continuous action spaces is an interesting direction for future work.
Assumption 2.
We can query the private expert policies, i.e., evaluate $\pi_{i}(a|s)$ for a given $i,s$ and $a$.
Again, this query access only comes up in computing the stability of the trajectory prefix, and can likely be relaxed to using behavior-cloned versions of the experts with a slightly worse utility bound.
Assumption 3.
Given $p_{\min}\in(0,1/|A|)$, define
$\Pi_{p_{\min}}=\{\pi\colon\inf_{s\in S}\min_{a\in A}\pi(a|s)\geq p_{\min}\},$
as the set of all stochastic policies with minimum action probability at least $p_{\min}$. Then all expert policies are in this set, that is, $\Pi\subseteq\Pi_{p_{\min}}$ and our privacy guarantee only considers policies in $\Pi_{p_{\min}}$.
This last assumption is purely technical and an artifact of our analysis. Without a guaranteed $p_{\min}$ the ratio of action probabilities between two neighboring databases can be infinite, causing a problem in our proofs. We look forward to removing this assumption in future work.
@@ 4 ALGORITHM
[…]
@@
2:Compute query count as
3:$count_{\tau}(\Pi)\leftarrow\sum_{\pi\in\Pi}\prod_{j=1}^{|\tau|}\pi(a_{j}|s_{j})$
4:if $count_{\tau}(\Pi)+Lap(\frac{4}{\varepsilon})>\hat{\theta}$ then
5:     Return $\top$
6:else
7:     Return $\perp$
@@
1:Expert dataset $D^{\Pi}$, expert policies $\Pi$, $\varepsilon_{1}$, $\delta_{1}$, min action-selection probability $p_{min}$, unstable query cutoff $T\leq N\cdot m$, trajectory length $L$
2:$\varepsilon^{\prime}\leftarrow\frac{\varepsilon_{1}}{\sqrt{32T\log(2/\delta_{1})}},\quad\delta^{\prime}\leftarrow\frac{\delta_{1}}{(2TL)}$
3:$c_{min}\leftarrow\frac{e^{\varepsilon^{\prime}}}{(e^{\varepsilon^{\prime}}-1)},\quad\theta\leftarrow\frac{c_{min}}{p_{min}},\quad D^{\Pi}_{stable}\leftarrow\{\}$
4:Reshuffle the order of $D^{\Pi}$
5:for $c=1,\dots,T$ do
6:     $\tau\leftarrow$ next trajectory from $D^{\Pi}$
7:     $\hat{\theta}\leftarrow\theta+\frac{4}{\varepsilon^{\prime}}\log(1/\delta^{\prime})+Lap(\frac{2}{\varepsilon^{\prime}})$
8:     for $i=1$ to $L$ do
9:         $r\leftarrow\mathcal{A}_{PQ}(\Pi,\tau_{1:i},\hat{\theta},\varepsilon^{\prime})$triangleright Algorithm 1
10:         if $r=\top$ then
11:              if $i=L$ then
12:                  $D^{\Pi}_{stable}\leftarrow D^{\Pi}_{stable}\cup\{\tau_{1:i}\}$
13:         else
14:              if $i\neq 1$ then
15:                  $D^{\Pi}_{stable}\leftarrow D^{\Pi}_{stable}\cup\{\tau_{1:i-1}\}$
16:              break
17:return $D^{\Pi}_{stable}$
Algorithm 2 uses Algorithm 1 as a subroutine, traversing up to $T$ trajectories as long as the prefix remains stable. The dataset $D^{\Pi}_{stable}$ returned by Algorithm 2 is composed of these stable trajectory prefixes. Any trajectories after the first $T$ for a suitably chosen parameter $T$ are not included in $D^{\Pi}_{stable}$ to respect the privacy budget. The discarded tails of each trajectory (including the trajectories that Algorithm 2 did not process) are collected in another dataset:
$D^{\Pi}_{unst}=\{\tau_{k+1:L}\mbox{ if }\tau_{1:k}\in D^{\Pi}_{stable}\mbox{ for $k\geq 1$ or }\tau,~{}~{}\tau:D^{\Pi}\}$
Consider any parameterized gradient-based offline RL algorithm $\mathcal{H}_{\phi}$. Training this algorithm consists of using $(s,a,r,s^{\prime})$ tuples collected in the offline dataset to update the parameters (by splitting the trajectories into into individual transitions). Having split the dataset into stable and unstable trajectories, we can now leverage the stability of $D^{\Pi}_{stable}$ to accumulate the gradient updates of $\mathcal{H}_{\phi}$ on transitions drawn from this dataset without any noise addition. The full algorithm, detailing the training using a mix of noisy gradient updates for transitions from $D^{\Pi}_{unst}$ and noiseless gradient updates for transitions from $D^{\Pi}_{stable}$ is given in Algorithm 3.
@@
1:Stable dataset $D^{\Pi}_{stable}$, Unstable dataset $D^{\Pi}_{unst}$, Sampling probability $p$, Gradient-based offline RL algorithm $\mathcal{H}_{\phi}$, $\varepsilon_{2},\delta_{2}$, $N$ training steps, $B$ batch size
@@
3:for $n=1$ to $N$ do
4:     $b\sim Bernoulli(p)$
5:     if $b=1$ then $\triangleright$ w.p. $p$
6:         Sample batch of transitions $\mathcal{B}\sim D^{\Pi}_{unst}$
7:         $\phi_{n+1}\leftarrow$ DP-SGD($\phi_{n},\mathcal{B},\varepsilon_{2},\delta_{2},N$)
8:     $\triangleright$ w.p. $1-p$
9:         Sample batch of transitions $\mathcal{B}\sim D^{\Pi}_{stable}$
10:         $\phi_{n+1}\leftarrow$ SGD($\phi_{n},\mathcal{B}$)
11:return $\mathcal{H}_{\phi_{N}}$
The sampling probability $p$ controls the expected fraction of timesteps sampled from the unstable dataset $D^{\Pi}_{unst}$. The DP-SGD updates of the form DP-SGD($\phi_{n},\mathcal{B},\varepsilon_{2},\delta_{2},N$) in Algorithm 3 refer to noisy gradient updates with additive noise scaled as to ensure $(\varepsilon_{2},\delta_{2})$ expert-level DP. To do so we add the modifications to the standard DP-SGD algorithm (Abadi et al.,, 2016):
  • To account for privacy amplification while sampling batches from the dataset, we use the use the fraction of experts from which we sample. Hence, subsampling coefficient $q$ for DP-SGD analysis is $q=p\cdot\frac{|B|}{m}$, where $m$ is the number of experts.
A detailed description of the modified expert-level DP-SGD, which we also employ as a baseline in our empirical work in Section 6 is provided in Appendix B. The privacy analysis for the overall pipeline combining Algorithms 2 and 3 is discussed in Section 5.
Sub-optimality guarantees for the learned policy: While we show in the following section that the Algorithm 3 provides expert-level privacy, we do not provide explicit bounds on the suboptimality of the returned policy. Typical offline RL literature shows that using pessimistic approaches, the learned policy is competitive with any other policy whose state-action distribution is well-covered by the offline data distribution (Xie et al.,, 2021; Cheng et al.,, 2022). However, our training distribution undergoes two changes from the apriori offline data distribution. First is that the stable prefixes might be distributionally rather different from the entire state-action distribution, and potentially emphasize the states and actions near the start of a trajectory, as shorter prefixes are more likely to be stable. Second, the gradient clipping in DP-SGD adds an additional bias. Understanding the effect of latter with arbitrary function approximation is beyond the scope of this paper. We do, however, study the effects of distributional biases from adding varying amounts of the stable prefixes in our experiments, and find that they are typically very beneficial, particularly in tasks where the optimal behaviour does not change significantly over time.
@@ 5 PRIVACY ANALYSIS
For the privacy analysis of our proposed algorithm, we adopt notations for the MDP described in Section 3.1 and problem setup described in Section 3.3. We assume for the benefit of the analysis, that all such trajectories are of the same length $L$ (number of actions taken before the episode ends). Practically this can be replaced by the length of the longest trajectory for a worst-case privacy analysis.
Algorithm 2 can be thought of as $T$ iterations of a combination of two algorithms $\mathcal{A}_{1},\mathcal{A}_{2}$:
  • $\mathcal{A}_{1}$ samples a random expert $\pi\sim\mathrm{Unif}(\Pi)$ and a trajectory $\tau\sim\pi$ and converts it into $L$ queries, $Q^{\tau}=\{count_{\tau_{1}}(.),...,count_{\tau_{L}}(.)\}$.
  • $\mathcal{A}_{2}$ is invoked on $\Pi$ and the query set $Q^{\tau}$, running an iteration of the Sparse Vector Technique (Dwork and Roth, (2014)). $\mathcal{A}_{2}$ treats $\Pi$ as the private database, $Q^{\tau}$ as the query sequence, and $\hat{\theta}$ as the threshold for the Sparse Vector Algorithm. $\mathcal{A}_{2}(\Pi,Q^{\tau})$ outputs $r\in\{\top,\bot\}^{k+1}$ such that $r[i]=\top\;\forall\;i\leq k$ and $r[k+1]=\bot$, where $k\leq L$ is the length of the stable prefix. Note that $r[i]=\mathcal{A}_{PQ}(\Pi,\tau_{1:i},\hat{\theta},\varepsilon^{\prime})$.
@@
For a fixed stream of queries $Q^{\tau}$, Algorithm $\mathcal{A}_{2}$ is $(\varepsilon^{\prime},0)$-DP since it just invokes one round of the Sparse Vector Algorithm (Dwork and Roth, (2014)). For a given trajectory $\tau$, $\mathcal{A}_{2}$ identifies the longest prefix which is stable, or in other words, has enough experts which are likely to sample this prefix. Note that, the privacy cost $\varepsilon^{\prime}$ for identifying this prefix is independent of the length of the trajectory.
@@
We now state two key lemmas for proving that $\mathcal{A}$ is differentially private. For a given trajectory prefix $\tau_{k}$ of length $k$, let $E_{\tau}$ denote the event that $\mathcal{A}_{1}$ samples any trajectory $\tau^{\prime}$ containing $\tau_{k}$ as a prefix; that is, $\tau^{\prime}_{1:k}=\tau_{k}$.
@@
This lemma states that if the probability of seeing a trajectory prefix $\tau_{k}$ across all experts is large enough, then the probability that $\mathcal{A}_{1}$ samples any trajectory with prefix $\tau_{k}$ is stable under changing one of the experts. The full proof is available in Appendix A.
@@
Informally, this lemma states that if the total probability of seeing a trajectory prefix $\tau_{k}$ across all experts is not large enough, then the probability that it is labelled as stable by Algorithm 1 is less than $\delta^{\prime}$. The full proof is described in Appendix A.
@@
Under Assumptions 1-3, $\mathcal{A}$ is $(2\varepsilon^{\prime},\frac{\delta_{1}}{2T})$-DP, where $\varepsilon^{\prime},\delta_{1}$ are defined in Algorithm 2.
@@
$\sum_{\tau\in\mathcal{T}}\max\{\Pr(\mathcal{A}(\Pi)=\tau)-e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau),0\}$
$\leq\frac{\delta_{1}}{2T}.$
The above can be proved using (4), and we defer the full proof to Appendix A.
@@
Fix a trajectory prefix $\tau_{k}$ and a pair of neighboring expert sets $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$. If $count_{\tau_{k}}(\Pi)<c_{min}$ then,
@@
$\Pr(\mathcal{A}(\Pi)=\tau_{k})$
$=\sum_{a\in A}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$
@@
(a) LunarLander
(b) Acrobot
(c) CartPole
(d) HIV Treatment
Figure 2: Expert return histograms for experts trained on heterogeneous LunarLander, Acrobot, CartPole and HIV Treatment environments. We use a mix of experts that shows a wide range of performance on the test environment. For HIV Treatment, the return values are normalised by $61e$ 6.
Let’s consider the case $k=L$. The main idea behind this proof is that if $count_{\tau_{k}}(\Pi)$ is small enough, then $\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau}_{k})=\gamma^{(L)})$ is less than $\delta^{\prime}$ (Lemma 5.2) and we get (5). On the other hand, if $count_{\tau_{k}}(\Pi)$ is greater than $c_{min}$, then $\Pr(E_{\tau_{k}}|\Pi)$ is bounded by Lemma 5.1. Using that $\mathcal{A}_{2}$ is always $\varepsilon^{\prime}$-DP when the queries are fixed we get (6). The case of $k<L$ is less straightforward, but can also be analyzed by considering two cases for each $a\in A$, $count_{\tau_{k}\cdot a}(\Pi)\geq c_{min}$ or $count_{\tau_{k}\cdot a}(\Pi)<c_{min}$. The full proof is present in Appendix A. ∎
Theorem 5.5.
Under Assumptions 1-3, Algorithm 2 is $(\varepsilon_{1},\delta_{1})$-DP.
@@
Algorithm 2 simply executes Algorithm $\mathcal{A}$, $T$ times. Since $\mathcal{A}$ is $(\varepsilon_{1}/\sqrt{8T\log(2/\delta_{1})},\delta_{1}/(2T))$-DP as shown in Theorem 5.3, using advanced composition (Corollary 3.21 in Dwork and Roth, (2014)), Algorithm 2 can be shown to be $(\varepsilon_{1},\delta_{1})$-DP. ∎
@@
Shown trivially by sequential composition of privacy guarantees for the release of $D^{\Pi}_{stable}$ and DP-SGD training on samples from $D^{\Pi}_{unst}$. ∎

### v2

@@ 3.3 Problem Setup
Let $\Pi$ denote a set of $m$ privacy-sensitive policies, corresponding to $m$ experts, $\Pi=\{\pi_{1},\dots,\pi_{m}\}$. $D^{\Pi}=D^{\pi_{1}}\cup\dots\cup D^{\pi_{m}}$ is the aggregated dataset generated by the interaction of the $m$ experts with the environment. Each $D^{\pi_{i}}$ is a set of $N$ trajectories $\tau=(s_{1},a_{2},s_{2},a_{2},\dots,s_{L},a_{L},s_{L+1})$ of length $L$ obtained by logging the interactions of policy $\pi_{i}$ with the MDP $\mathcal{M}=\langle S,A,r,P,\rho_{0},\gamma\rangle$, where $s_{1}\sim\rho_{0}$, $a_{h}\sim\pi_{i}(\cdot|s_{h})$ and $s_{h+1}\sim P(\cdot|s_{h},a_{h})$. We denote by $|\tau|$ the length of the trajectory as the number of full state-action pairs it consists of. Further, we use $\tau_{h:h^{\prime}}=(s_{h},a_{h},\dots s_{h^{\prime}},a_{h^{\prime}},s_{h^{\prime}+1})$ , where $h\leq h^{\prime}$, to denote the sub-trajectory of $\tau$ starting from the $h$-th timestep until the $h^{\prime}$-th, including the trailing next state $s_{h^{\prime}+1}$.
Note on Trajectory Notation: As in Definition 3.1, trajectories are classically defined as a series of $(s,a,r,s^{\prime})$ tuples, where the next state $s^{\prime}$ of one transition is the state $s$ of the next transition. Since the rewards associated with transitions are not used in the privacy analysis or in the filtering stage of our algorithm, we simplify the classical notation for our analysis and proofs as described above.
@@
The goal is to learn a policy using an offline RL algorithm with expert-level DP guarantees on $D^{\Pi}$, thus ensuring that the policy learnt in this manner is not much different from a policy learnt using $D^{\Pi^{\prime}}$.
We make the following mild assumptions in the design of our solution.
Assumption 3.4.
@@
This is a fairly mild assumption satisfied in many offline RL settings. This arises while identifying stable trajectory prefixes for our algorithm, and we leave an extension to continuous action spaces to future work.
Assumption 3.5.
We can query the sensitive expert policies, i.e., evaluate $\pi_{i}(a|s)$ for a given $i,s$ and $a$.
Similar assumptions are reasonable to make in setups involving learning a student model from an ensemble of privacy-sensitive teacher models, where access to the teacher models are assumed (e.g. papernot2022semi). The privacy-sensitive experts in our setting are analogous to these privacy-sensitive teacher models. Again, query access only comes up in computing the stability of the trajectory prefix, and can likely be relaxed to using behavior-cloned versions of the experts with a slightly worse utility bound.
Assumption 3.6.
Let $\Pi_{s}$ be the class of expert policies considered by our algorithm. Hence, $\Pi\subseteq\Pi_{s}$. We assume a minimum action probability ($p_{min}$), defined as, $p_{\min}\leq\min_{\pi\in\Pi_{s}}\inf_{s\in S}\min_{a\in A}\pi(a|s)$
This value is used in the data filtering stage (first stage) of our algorithm. Note that we can always take $p_{min}=0$. However, our algorithm is more meaningful when $p_{min}>0$. Examples of common policy classes with $p_{min}>0$ include $\varepsilon$-greedy policies and softmax policies with bounded parameters.
Existing work in DP-RL relies on strict assumptions like tabular settings or linear function approximation, limiting real-world applicability. Our approach removes these constraints and extends to any state-of-the-art gradient-based offline RL algorithm with general function approximators and continuous state spaces.
@@ 4 ALGORITHM
For a given trajectory prefix $\tau$ , we define the total probability of $\tau$ being generated by all the experts as: $count_{\tau}(\Pi)=\sum_{i=1}^{m}\left[\prod_{j=1}^{L_{\tau}}\pi_{i}(a_{j}|s_{j})\right]$  [^2] The full probability of being generated by the given set of experts, is actually proportional to the count function multiplied by the state-transition probabilities (which do not change for neighbouring experts)., where $L_{\tau}=|\tau|$, is the length of trajectory $\tau$. Algorithm 1 deems $\tau$ stable if $count_{\tau}(\pi)$ exceeds a carefully chosen threshold, implying that enough experts are likely to generate $\tau$. These stable prefixes may be used in training an offline RL algorithm without any noise addition. We need Assumption 3.4 for this step. Given a trajectory, the count function gives the sum of probabilities of each expert traversing it. We use $count_{\tau}(\Pi)$ to check if enough experts can generate this trajectory.
@@
2:Compute expected experts: $count_{\tau}(\Pi)\leftarrow\sum_{\pi\in\Pi}\prod_{j=1}^{|\tau|}\pi(a_{j}|s_{j})$
3:if $count_{\tau}(\Pi)+Lap(\frac{4}{\varepsilon})>\hat{\theta}$ then return $\top$
4:else return $\perp$
@@
1:Expert dataset $D^{\Pi}$, expert policies $\Pi$, $\varepsilon_{1}$, $\delta_{1}$, min action-selection probability $p_{min}$, unstable query cutoff $T\leq N\cdot m$
[…]
5:Reshuffle $D^{\Pi}$
6:for $c=1,\dots,T$ do
7:  $\tau\leftarrow$ next trajectory from $D^{\Pi}$
8:  $\hat{\theta}\leftarrow\theta+\frac{4}{\varepsilon^{\prime}}\log(1/\delta^{\prime})+Lap(\frac{2}{\varepsilon^{\prime}})$
9:  for $i=1$ to $L$ do
10:   $r\leftarrow\mathcal{A}_{PQ}(\Pi,\tau_{1:i},\hat{\theta},\varepsilon^{\prime})$
11:   if $r=\top$ and $i=|\tau|$ then $D^{\Pi}_{stable}\leftarrow D^{\Pi}_{stable}\cup\{\tau_{1:i}\}$
12:   else if $r=\perp$ then
13:     if $i>1$ then $D^{\Pi}_{stable}\leftarrow D^{\Pi}_{stable}\cup\{\tau_{1:i-1}\}$
14:     break
15:return $D^{\Pi}_{stable}$
Algorithm 2 uses Algorithm 1 as a subroutine, traversing up to $T$ trajectories to find stable prefixes. The dataset $D^{\Pi}_{stable}$ returned by Algorithm 2 is composed of these stable trajectory prefixes. The discarded tails of each trajectory (including several full trajectories that were discarded) are collected in another dataset:
$D^{\Pi}_{unst}=\{\tau_{k+1:|\tau|}\mbox{ if }\tau_{1:k}\in D^{\Pi}_{stable}\mbox{ for $k\geq 1$ or }\tau,~~\tau:D^{\Pi}\}$
Consider any parameterized gradient-based offline RL algorithm $\mathcal{H}_{\phi}$, where $\phi$ denotes the parameters. Training this algorithm consists of using $(s,a,r,s^{\prime})$ tuples collected in the offline dataset to update the parameters (by splitting the trajectories into into individual transitions). Having split the dataset into stable and unstable trajectories, we can now leverage the stability of $D^{\Pi}_{stable}$ to accumulate the gradient updates of $\mathcal{H}_{\phi}$ on transitions drawn from this dataset without any noise addition. The full algorithm, detailing the training using a mix of noisy gradient updates for transitions from $D^{\Pi}_{unst}$ and noiseless gradient updates for transitions from $D^{\Pi}_{stable}$ is given in Algorithm 3.
The sampling probability $p$ controls the expected fraction of batches sampled from the unstable dataset $D^{\Pi}_{unst}$. The DP-SGD updates of the form DP-SGD($\phi_{n},\mathcal{B},\varepsilon_{2},\delta_{2},N$) in Algorithm 3 refer to noisy gradient updates with additive noise scaled as to ensure $(\varepsilon_{2},\delta_{2})$ expert-level DP. To do so we add the modifications to the standard DP-SGD algorithm (abadi2016dp):
  • To account for privacy amplification while sampling batches from the dataset, we use the use the fraction of experts from which we sample. Hence, the subsampling coefficient $q$ for DP-SGD analysis is $q=p\cdot\frac{|B|}{m}$, where $m$ is the number of experts.
We detail the modified expert-level DP-SGD, used as a baseline in Section 6, in Appendix B. See Section 5 for the privacy analysis for the overall pipeline combining Algorithms 2 and 3.
@@
1:Stable dataset $D^{\Pi}_{stable}$, Unstable dataset $D^{\Pi}_{unst}$, Sampling probability $p$, Gradient-based offline RL algorithm $\mathcal{H}_{\phi}$, $\varepsilon_{2},\delta_{2}$, NN training steps, $B$ batch size
@@
3:for $n=1$ to NN do
4:  $b\sim\text{Bernoulli}(p)$
5:  if $b=1$ then $\phi_{n+1}\leftarrow\text{DP-SGD}(\phi_{n},\mathcal{B},\varepsilon_{2},\delta_{2},N)$, where we sample batch $\mathcal{B}\sim D^{\Pi}_{unst}$
6:  else if $b=0$ then $\phi_{n+1}\leftarrow\text{SGD}(\phi_{n},\mathcal{B})$, where we sample batch $\mathcal{B}\sim D^{\Pi}_{stable}$
7:return $\mathcal{H}_{\phi_{N}}$
Relationship with User-level DP: The relationship between trajectory-level and expert-level DP is analogous to the notions of example-level and user-level privacy in existing DP literature (levy2021user; zhou2022locallydp; zhao2024huber). However, we retain naming conventions from existing RL literature.
We note that the expert-level DP-SGD (discussed above and in Appendix B) update step in Line 6 of Algorithm 3 is compatible with other user-level DP algorithms that can be used for training on the unstable data instead. Most practical approaches for achieving user-level DP are based on clipping user-level gradients to bound user contributions (levy2021user; bassily2023userlevel; liu2024userlevel; de2022dpimage). These approaches have been empirically tested to be successful on large scale datasets (charles2024finetuning). Other approaches for achieving user-level DP are mostly theoretical in nature and come with restrictions (such as on the function class) that limit their practical usage. The baseline used for comparison in our paper (described in Appendix B) is a representative of this class of approaches, and a natural way to tackle our problem statement in such a way that we generalize to any function approximators and continuous, high-dimensional state spaces. We demonstrate that our method outperforms this baseline by identifying a subset of the data where noisy training is not required, before running the same baseline on the rest of the data.
Sub-Optimality Guarantees for the Learned Policy: While we show in the following section that the Algorithm 3 provides expert-level privacy, we do not provide explicit bounds on the suboptimality of the returned policy. Typical offline RL literature shows that using pessimistic approaches, the learned policy is competitive with any other policy whose state-action distribution is well-covered by the offline data distribution (xie2021bellman; cheng2022adversarially). However, our training distribution undergoes two changes from the a priori offline data distribution. First is that the stable prefixes might be distributionally rather different from the entire state-action distribution, and potentially emphasize the states and actions near the start of a trajectory, as shorter prefixes are more likely to be stable. Second, the gradient clipping in DP-SGD adds an additional bias. Understanding the effect of latter with arbitrary function approximation is beyond the scope of this paper. We do, however, study the effects of distributional biases from adding varying amounts of the stable prefixes in our experiments, and find that they are typically very beneficial, particularly in tasks where the optimal behaviour does not change significantly over time.
@@ 5 PRIVACY ANALYSIS
We note that all the trajectories may be of various lengths. However, as we shall see in the analysis, the privacy cost in the $\delta$ parameter scales linearly with the length of the trajectory, i.e., $\delta_{1}\propto L\delta^{\prime}$ (Line 2 in Algorithm 2), where $\delta^{\prime}$ is the per-timestep cost and $\delta_{1}$ is the full cost of the trajectory. By setting $L$ to be the maximum length across all trajectories, we carry out a worst-case analysis of the privacy leakage. For a fixed per-timestep privacy budget, we overestimate the total privacy cost by safely considering the worst-case (maximum) trajectory lengths. In other words, for the benefit of the analysis, we assume that all trajectories have the same (maximal) length $L=\underset{\tau\in D^{\Pi}}{\text{max}}\;|\tau|$.
Algorithm 2 can be thought of as $T$ iterations of a combination of two algorithms $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$:
  • $\mathcal{A}_{1}$ samples a random expert $\pi\sim\mathrm{Unif}(\Pi)$ and a trajectory $\tau\sim D^{\pi}$ and converts it into $|\tau|$ queries, $Q^{\tau}=\{count_{\tau_{1}}(.),\dots,count_{\tau_{|\tau|}}(.)\}$, where $\tau_{i}=\tau_{1:i}$ is an $i$-length prefix of the trajectory $\tau$.
  • $\mathcal{A}_{2}$ is invoked on $\Pi$ and the query set $Q^{\tau}$, running an iteration of the Sparse-Vector Technique (dwork2014dp) (SVT). $\mathcal{A}_{2}$ treats $\Pi$ as the privacy-sensitive database, $Q^{\tau}$ as the query sequence, and $\hat{\theta}$ as the threshold for SVT. $\mathcal{A}_{2}(\Pi,Q^{\tau})$ outputs $r\in\{\top,\bot\}^{k+1}$ such that $r[i]=\top\;\forall\;i\leq k$ and $r[k+1]=\bot$, where $k\leq L$ is the length of the stable prefix. Note that $r[i]=\mathcal{A}_{PQ}(\Pi,\tau_{1:i},\hat{\theta},\varepsilon^{\prime})$.
@@
For a fixed stream of queries $Q^{\tau}$, Algorithm $\mathcal{A}_{2}$ is $(\varepsilon^{\prime},0)$-DP since it just invokes one round of the Sparse Vector Algorithm (dwork2014dp). Note that the trajectories are just used for creating the queries for SVT, and the set of expert policies $\Pi$ is the privacy database. For a given trajectory $\tau$, $\mathcal{A}_{2}$ identifies the longest prefix which is stable, or in other words, has enough experts which are likely to sample this prefix. Note that, the privacy cost $\varepsilon^{\prime}$ for identifying this prefix is independent of the length of the trajectory.
@@
We now state two key lemmas for proving that $\mathcal{A}$ is differentially private. For a given trajectory prefix $\tau_{k}$ of length $k$, let $E_{\tau}$ denote the event that $\mathcal{A}_{1}$ samples any trajectory $\tau^{\prime}$ containing $\tau_{k}$ as a prefix, i.e., $\tau^{\prime}_{1:k}=\tau_{k}$.
@@
If the probability of seeing a trajectory prefix $\tau_{k}$ across all experts is large enough, then the probability that $\mathcal{A}_{1}$ samples any trajectory with prefix $\tau_{k}$ is stable when one expert changes. See Appendix A for proof.
@@
Informally, if the total probability of seeing a trajectory prefix $\tau_{k}$ across all experts is not large enough, then the probability that it is labelled as stable by Algorithm 1 is less than $\delta^{\prime}$. See Appendix A for proof.
@@
Under Assumption 3.4, $\mathcal{A}$ is $(2\varepsilon^{\prime},\frac{\delta_{1}}{2T})$-DP, where $\varepsilon^{\prime},\delta_{1}$ are defined in Algorithm 2.
@@
$\sum_{\tau\in\mathcal{T}}\max\{\Pr(\mathcal{A}(\Pi)=\tau)-e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau),0\}\leq\frac{\delta_{1}}{2T}.$
The above can be proved using equation 4, and we defer the full proof to Appendix A.
@@
For a trajectory prefix $\tau_{k}$ and neighboring expert sets $\Pi,\Pi^{\prime}$, if $count_{\tau_{k}}(\Pi)<c_{min}$ then,
@@
$\Pr(\mathcal{A}(\Pi)=\tau_{k})=\sum_{a\in A}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$
@@
Consider the case $k=L$. If $count_{\tau_{k}}(\Pi)$ is small enough, then $\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau}_{k})=\gamma^{(L)})$ is less than $\delta^{\prime}$ (Lemma 5.2) and we get equation 5. On the other hand, if $count_{\tau_{k}}(\Pi)$ is greater than $c_{min}$, then $\Pr(E_{\tau_{k}}|\Pi)$ is bounded by Lemma 5.1. Using that $\mathcal{A}_{2}$ is always $\varepsilon^{\prime}$-DP when the queries are fixed we get equation 6. The case of $k<L$ is less straightforward, but can also be analyzed by considering two cases for each $a\in A$, $count_{\tau_{k}\cdot a}(\Pi)\geq c_{min}$ or $count_{\tau_{k}\cdot a}(\Pi)<c_{min}$. See Appendix A for the full proof. ∎
Theorem $5.$
Under Assumption 3.4, Algorithm 2 is $(\varepsilon_{1},\delta_{1})$-DP.
@@
Algorithm 2 executes $\mathcal{A}$ with $(\varepsilon_{1}/\sqrt{8T\log(2/\delta_{1})},\delta_{1}/(2T))$-DP (see Theorem 5.3), $T$ times. Using advanced composition (Corollary 3.21 in dwork2014dp) we get that Algorithm 2 is $(\varepsilon_{1},\delta_{1})$-DP. ∎
@@
Follows from sequential composition of DP guarantees for the release of $D^{\Pi}_{stable}$ and DP-SGD training on samples from $D^{\Pi}_{unst}$. ∎
Note on Novelty: We present a novel, useful and non-trivial extension of the Sparse Vector Technique (SVT) to preserve expert-level DP in Offline RL. The stream of queries in vanilla SVT, (Algorithm 1, Section 3.6, dwork2014dp), is not dependent on the private database. However, in our setup, the stream of queries is a function of the set of the experts (which is the private database to be protected), and directly applying SVT does not ensure privacy. However, we show that despite the query stream being a function of the private database, SVT can work if the threshold (line 6 in Algorithm 2) is chosen in a certain way. We use the probability distribution of the query stream derived from the experts’ policies to prove this.